OpenStudy (anonymous):

x^8 y^12 / x^4 y^4 x=2 y=-1

Directrix (directrix):

(x^8 y^12) / (x^4 y^4) = x ^ (8-4) * y ^ (12-4) = x^4 / y ^ 8 Check this work for accuracy. Then subsitute 2 for x and -1 for y.

OpenStudy (anonymous):

yes

Directrix (directrix):

Then subsitute 2 for x and -1 for y.

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

If you simplify the equation you get (x^4)*(y^8) so then you substitute 2 for x and -1 for y